# Beyond Seen Data: Improving KBQA Generalization Through Schema-Guided Logical Form Generation

### v2

@@ 4.1 Relation and Entity Retrieval
For relation retrieval, we follow the schema retrieval model of TIARA Shu et al. (2022), as it has high accuracy. We extract a set $R_{q}$ of top-$k_{R}$ (system parameter) relations with the highest semantic similarity to $q$. This is done by a BERT-based cross-encoder that learns the semantic similarity $\text{sim}(q,r)$ between $q$ and a relation $r\in\mathcal{R}$:
@@
Relation-Guided Entity Mention Detection
Given $R_{q}$, we propose a relation-guided logical form sketch parser to parse $q$ into a logical form sketch $s$. Entity mentions in $q$ are extracted from $s$.
@@
Relation-Guided Candidate Entity Retrieval
We follow previous studies Gu et al. (2021); Shu et al. (2022); Faldu et al. (2024); Luo et al. (2024) and use an entity name dictionary FACC1 Gabrilovich et al. (2013) to map extracted entity mentions to entities (i.e., their IDs in KB), although other retrieval models can be used. Since different entities may share the same name, the entity mentions may be mapped to many entities. For pruning, existing studies use popularity scores associated to entities Shu et al. (2022); Ye et al. (2022).
To improve the recall of candidate entity retrieval, we propose a combined pruning strategy based on both popularity and relation connectivity. As Figure 4 shows, we first select the top-$k_{E1}$ (system parameter) entities for each entity mention based on popularity and then extract $k_{E2}$ (system parameter) entities from the remaining candidates that are connected to the retrieved relations $R_{q}$. Together, these form the candidate entity set $E_{c}$.
Figure 4: Candidate entity retrieval for the mention ‘aloha’. The candidate entity in red is the ground-truth.
@@ 4.2 Schema-Guided Logical Form Generation
Before being fed into the model, each relation and entity is augmented with its schema context (i.e., class information) to help the model to learn their connections and generalize to unseen entities, relations, or their compositions. The context of a relation $r$ is described by concatenating the relation’s domain class $c_{d}$ and range class $c_{r}$, formatted as “[D] $c_{d}$ [N] $r$ [R] $c_{r}$”. For an entity $e$, its context is described by its ID (“$id_{e}$”), name (“$name_{e}$”), and the intersection of its set of classes $C_{e}$ and the set of all domain and range classes $C_{R}$ of all relations in $R_{q}$, formatted as “[ID] $id_{e}$ [N] $name_{e}$ [C] class($C_{e}\cap C_{R}$)”.
@@
$\mathcal{L}_{generator}=-\sum_{t=1}^{n}\log p\left(l_{t}\mid l_{<t},q,K_{q}\right),$ (2)
where $l$ denotes a logical form of $n$ tokens and $l_{t}$ is its $t$-th token, and $K_{q}$ is the retrieved knowledge (i.e., relations and entities with contexts) for $q$. At inference, the model runs beam search to generate top-$k_{O}$ logical forms – the executable one with the highest score is selected as the output. See Appendix B for a prompt example used for inference.
It is possible that no generated logical forms are executable. In this case, we fall back to following Shu et al. (2022) and Ye et al. (2022) and retrieve candidate logical forms in two stages: enumeration and ranking. During enumeration, we search the KB by traversing paths starting from the retrieved entities. Due to the exponential growth in the number of candidate paths with each hop, we start from the top-1 entity for each mention and searches its neighborhood for up to two hops. The paths retrieved are converted into logical forms. During ranking, a BERT-based ranker scores $q$ and each enumerated logical form $l$ (like Eq. 1). We train the ranker using a contrastive objective:
@@
where $l^{*}$ is the ground-truth logical form and $C_{l}$ is the set of enumerated logical forms. We run the ranked logical forms from the top and return the first executable one.
@@ 5 Experiments
We run experiments to answer: Q1: How does SG-KBQA compare with SOTA models in their accuracy for the KBQA task? Q2: How do model components impact the accuracy of SG-KBQA? Q3: How do our techniques generalize to other KBQA models?
@@ 5.1 Experimental Setup
Following SOTA competitors Shu et al. (2022); Gu et al. (2023); Zhang et al. (2023), we use two benchmark datasets built upon Freebase.
@@
GrailQA Gu et al. (2021) is a dataset for evaluating the generalization capability of KBQA models. It contains 64,331 questions with annotated target S-expressions, including complex questions requiring up to 4-hop reasoning over Freebase, with aggregation functions including comparatives, superlatives, and counting. The dataset comes with training (70%), validation (10%), and test (20%, hidden and only known by the leaderboard organizers) sets. In the validation and the test sets, 50% of the questions include KB elements that are unseen in the training set (zero-shot generalization tests), 25% consist of unseen compositions of KB elements seen in the training set (compositional generalization tests), and the remaining 25% are randomly sampled from the training set (I.I.D. tests).
@@
WebQuestionsSP (WebQSPYih et al. (2016) is a dataset for the I.I.D. setting. While our focus is on non-I.I.D. settings, we include results on this dataset to show the general applicability of SG-KBQA. WebQSP contains 4,937 questions. More details of WebQSP are included in Appendix C.
@@
SP-based (SFT) RnG-KBQA (ACL 2021) 68.8 74.4 86.2 89.0 63.8 71.2 63.0 69.2
TIARA (EMNLP 2022) 73.0 78.5 87.8 90.6 69.2 76.5 68.0 73.9
Decaf (ICLR 2023) 68.4 78.7 84.8 89.9 73.4 81.8 58.6 72.3
Pangu (T5-3B) (ACL 2023) 75.4 81.7 84.4 88.8 74.6 81.5 71.6 78.5
FC-KBQA (ACL 2023) 73.2 78.7 88.5 91.2 70.0 76.7 67.6 74.0
TIARA+GAIN (EACL 2024) 76.3 81.5 88.5 91.2 73.7 80.0 71.8 77.8
RetinaQA (ACL 2024) 74.1 79.5 - - 71.9 78.9 68.8 74.7
SP-based (Few-shot) KB-Binder (6)-R (ACL 2023) 53.2 58.5 72.5 77.4 51.8 58.3 45.0 49.9
Pangu (Codex) (ACL 2023) 56.4 65.0 67.5 73.7 58.2 64.9 50.7 61.1
FlexKBQA (AAAI 2024) 62.8 69.4 71.3 75.8 59.1 65.4 60.6 68.3
Ours (SFT) SG-KBQA 79.1 84.4 88.6 91.6 77.9 85.1 75.4 80.8
  - Improvement +3.6% +3.3% +0.1% +0.4% +4.4% +4.0% +5.0% +2.9%
Table 1: Hidden test results (%) on GrailQA (best results are in boldface; best baseline results are underlined; “SFT” means supervised fine-tuning; “few-shot” means few-show in-context learning).
Competitors
On GrailQA, we compare with models that top the leaderboard [^1] https://dki-lab.github.io/GrailQA/, including RnG-KBQA Ye et al. (2022), TIARA Shu et al. (2022), DecAF Yu et al. (2023), Pangu (previous SOTA as of 15th February, 2025) Gu et al. (2023), FC-KBQA Zhang et al. (2023), TIARA+GAIN Shu and Yu (2024), and RetinaQA Faldu et al. (2024). We also compare with few-shot LLM (training-free) methods: KB-BINDER (6)-R Li et al. (2023), Pangu  Gu et al. (2023), and FlexKBQA Li et al. (2024). These models are SP-based. On the non-I.I.D. GrailQA, IR-based methods are uncompetitive and excluded.
On WebQSP, we compare with IR-based models SR+NSM Zhang et al. (2022), UNIKGQA Jiang et al. (2023), and EPR+NSM Ding et al. (2024), plus SP-based models ChatKBQA (SOTA) Luo et al. (2024) and TFS-KBQA (SOTA) Wang and Qin (2024), both of which use a fine-tuned LLM to generate logical forms. We also compare with TIARA, Pangu, and FC-KBQA as above, which represent SOTA models using pre-trained language models (PLMs). Appendix D details these models. The baseline results are collected from their papers or the GrailQA leaderboard (if available).
Implementation Details
All our experiments are run on a machine with an NVDIA A100 GPU and 120 GB of RAM. We fine-tuned three bert-base-uncased models for a maximum of three epochs each, for relation retrieval, entity ranking, and fallback logical form ranking. For each dataset, a T5-base model is fine-tuned for 5 epochs as our logical form sketch parser. Finally, we fine-tune a LLaMA3.1-8B with LoRA Hu et al. (2022a) for 5 epochs on GrailQA and 20 epochs on WebQSP to serve as the logical form generator. Our system parameters have been chosen empirically, and a parameter study is provided in Appendix H. More implementation details are in Appendix E.
Evaluation Metrics
On GrailQA, we report the exact match (EM) and F1 scores, following the leaderboard. EM counts the percentage of test samples where the model generated logical form (an S-expression) that is semantically equivalent to the ground truth. F1 measures the answer set correctness, i.e., the F1 score of each answer set, average over all test samples. On WebQSP, we report the F1 score as there are no ground-truth S-expressions.
Model F1
IR-based SR+NSM (ACL 2022) 69.5
UniKGQA (ICLR 2023) 75.1
EPR+NSM (WWW 2024) 71.2
SP-based (SFT) TIARA (EMNLP 2022) 76.7
Pangu (T5-3B, ACL 2023) 79.6
FC-KBQA (ACL 2023) 76.9
ChatKBQA (ACL 2024) 79.8
TFS-KBQA (LREC-COLING 2024) 79.9
SP-based (Few-shot) KB-Binder (6)-R (ACL 2023) 53.2
Pangu (Codex) (ACL 2023) 54.5
FlexKBQA (AAAI 2024) 60.6
Ours (SFT) SG-KBQA 80.3
     - Improvement +0.5%
Table 2: Test results (%) on WebQSP (I.I.D.).
5.2 Overall Results (Q1)
Tables 1 and 2 show the overall comparison of SG-KBQA with the baseline models for GrailQA and WebQSP, respectively. SG-KBQA shows the best results across both datasets.
Results on GrailQA
On the overall hidden test set of GrailQA, SG-KBQA outperforms the best baseline Pangu by 4.9% and 3.3% in the EM and F1 scores, respectively. Under the compositional and zero-shot generalization settings (both are non-I.I.D.), similar performance gaps are observed, i.e., 4.0% and 2.9% in F1 compared to the best baseline models, respectively. This validates that SG-KBQA can extract relations and entities more accurately from the input question, even when these are unseen in the training set, and it creates more accurate logical forms to answer the questions.
The fine-tuned baseline models do not use relation semantics to enhance entity retrieval, and they either omit the class contexts in logical form generation or use these classes separately for retrieval. As such, they do not generalize as well in the non-I.I.D. settings. The few-shot LLM-based competitors are generally not very competitive, especially under the non-I.I.D. settings. This suggests that the current generation of LLMs are unable to infer from a few input demonstrations the process of logical form generation from user questions. Fine-tuning is still required.
Results on WebQSP
On WebQSP, which has an I.I.D. test set, the performance gap of the different models are closer. Even in this case, SG-KBQA still performs the best, showing its general applicability. Comparing with TFS-KBQA (SOTA) and ChatKBQA, SG-KBQA improves the F1 score by 0.5%. Among IR-based methods, UniKGQA (SOTA) still performs substantially worse compared to SG-KBQA. The lower performance of IR-based methods is consistent with existing results Gu et al. (2022).

### v3

@@ 4.1 Relation Retrieval
For relation retrieval, we follow TIARA Shu et al. (2022) for its high accuracy. We extract a set $R_{q}$ of top-$k_{R}$ (system parameter) relations with the highest semantic similarity to $q$. This is done by a BERT-based cross-encoder to learn the semantic similarity between $q$ and a relation $r\in\mathcal{R}$:
@@
4.2 Schema-guided Entity Retrieval
Entity Mention Detection
Given $R_{q}$, we propose a schema-guided logical form sketch parser to parse $q$ into a logical form sketch $s$. Entity mentions in $q$ are extracted from $s$.
@@ 4.2 Schema-guided Entity Retrieval
Candidate Entity Retrieval
We follow previous studies Faldu et al. (2024); Luo et al. (2024a); Shu et al. (2022) and use an entity name dictionary FACC1 Gabrilovich et al. (2013) to map extracted entity mentions to entities (i.e., their IDs in KB), although other retrieval models can be used. Since different entities may share the same name, the entity mentions may be mapped to many entities. For pruning, existing studies use popularity scores of the entities Shu et al. (2022); Ye et al. (2022).
To improve the recall, we propose a combined pruning strategy based on both popularity and relations. As Figure 3 shows, we first select the top-$k_{E1}$ (system parameter) entities for each mention based on popularity and then extract $k_{E2}$ (system parameter) entities from the remaining candidates that are connected to the retrieved relations $R_{q}$. Together, these form the candidate entity set $E_{c}$.
Figure 3: Candidate entity retrieval for ‘Harry Potter’. The candidate entity in red is the ground-truth.
@@ 4.3 Schema-Guided Logical Form Generation
Before being fed into the model, each relation and entity is augmented with its class information to help the model learn their connections and generalize to unseen entities, relations, or their compositions. The context of a relation $r$ is described by concatenating its domain class $c_{d}$ and range class $c_{r}$, formatted as “[D] $c_{d}$ [N] $r$ [R] $c_{r}$”. For an entity $e$, its context is described by its ID ($id_{e}$”), name ($name_{e}$”), and the intersection between its set of classes $C_{e}$ and the set of all domain and range classes $C_{R}$ of all relations in $R_{q}$, formatted as “[ID] $id_{e}$ [N] $name_{e}$ [C] class($C_{e}\cap C_{R}$)”.
@@
$\mathcal{L}_{\text{generator}}=-\sum_{t=1}^{n}\log p\left(l_{t}\mid l_{<t},q,K_{q}\right),$ (2)
where $l$ denotes a logical form of $n$ tokens and $l_{t}$ is its $t$-th token, and $K_{q}$ is the retrieved knowledge (i.e., relations and entities with class contexts) for $q$. At inference, the model runs beam search to generate top-$k_{O}$ logical forms – the executable one with the highest score is the output. See Appendix B for a prompt example used for inference.
It is possible that no generated logical forms are executable. In this case, we fall back to following Shu et al. (2022) and Ye et al. (2022) and retrieve candidate logical forms in two stages: enumeration and ranking. During enumeration, we traverse the KB starting from the retrieved entities. Due to the exponential growth in the candidate paths with each hop, we start from the top-1 entity for each mention and examine its neighborhood for up to two hops. The paths are converted into logical forms. During ranking, a BERT-based ranker scores $q$ and each logical form $l$ (like Eq. 1). We train the ranker using a contrastive objective:
@@
where $l^{*}$ is the ground-truth logical form and $C_{l}$ is the set of enumerated logical forms. The top-ranked, executable logical form is returned.
Overall I.I.D. Compositional Zero-shot
Model EM F1 EM F1 EM F1 EM F1
SP-based (SFT) RnG-KBQA 68.8 74.4 86.2 89.0 63.8 71.2 63.0 69.2
TIARA 73.0 78.5 87.8 90.6 69.2 76.5 68.0 73.9
Decaf 68.4 78.7 84.8 89.9 73.4 81.8 58.6 72.3
Pangu (T5-3B) 75.4 81.7 84.4 88.8 74.6 81.5 71.6 78.5
FC-KBQA 73.2 78.7 88.5 91.2 70.0 76.7 67.6 74.0
TIARA+GAIN 76.3 81.5 88.5 91.2 73.7 80.0 71.8 77.8
RetinaQA 74.1 79.5 - - 71.9 78.9 68.8 74.7
SP-based (Few-shot) KB-Binder (6)-R 53.2 58.5 72.5 77.4 51.8 58.3 45.0 49.9
Pangu (Codex) 56.4 65.0 67.5 73.7 58.2 64.9 50.7 61.1
FlexKBQA 62.8 69.4 71.3 75.8 59.1 65.4 60.6 68.3
Ours (SFT) SG-KBQA 79.1 84.4 88.6 91.6 77.9 85.1 75.4 80.8
  - Improvement +3.6% +3.3% +0.1% +0.4% +4.4% +4.0% +5.0% +2.9%
Table 1: Hidden test results (%) on GrailQA (best results are in boldface; best baseline results are underlined; “SFT” means supervised fine-tuning; “few-shot” means few-shot in-context learning).
@@ 5 Experiments
We run experiments to answer: Q1: How does SG-KBQA improve generalizability compared with SOTA models? Q2: How do model components contribute to generalizability? Q3: How can our techniques enhance existing models?
@@ 5.1 Experimental Setup
Following SOTA competitors (Shu et al., 2022; Gu et al., 2023; Zhang et al., 2023), we use three benchmark datasets built upon Freebase.
@@
GrailQA (Gu et al., 2021) is a dataset for evaluating the generalizability of KBQA models. It has 64,331 questions with target S-expressions, including complex questions requiring up to 4-hop reasoning over the KG. The dataset comes with training (70%), validation (10%), and test (20%, hidden and only known by the leaderboard organizers) sets. In the validation and the test sets, 50% of the questions include KB elements that are unseen in the training set (zero-shot generalization tests), 25% consist of unseen compositions of KB elements seen in the training set (compositional generalization tests), and the remaining 25% are randomly sampled from the training set (I.I.D. tests).
@@
WebQuestionsSP (WebQSP(Yih et al., 2016) is a dataset for the I.I.D. setting. While our focus is on non-I.I.D. settings, we include results on this dataset to show the general applicability of SG-KBQA. WebQSP contains 4,937 questions collected from Google query logs, including 3,098 questions for training and 1,639 for testing, each annotated with a target SPARQL query. We follow GMT-KBQA Hu et al. (2022b) and TIARA Shu et al. (2022) to separate 200 questions from the training questions to form the validation set.
ComplexWebQuestions (CWQ(Talmor and Berant, ) is a challenging I.I.D. dataset for testing complex KBQA. It consists of 34,689 questions requiring up to 4-hop reasoning, constructed by extending the WebQSP dataset to questions with higher-hop complexity. For our experiments, we adopt the official split of CWQ, where 80% of the questions are allocated for training, 10% for validation, and the remaining 10% for testing.
Model WebQSP CWQ
IR-based
SR+NSM 69.5 47.1
UniKGQA 75.1 49.4
EPR+NSM 71.2 61.2
LLM-based
Pangu (ACL 2023) 79.6 -
RoG 70.8 56.2
ChatKBQA 79.8 77.8
TFS-KBQA 79.9 63.6
LLM-based
(Few-shot)
KB-Binder (6)-R 53.2 -
Pangu (Codex) 54.5 -
FlexKBQA 60.6 -
ToG (GPT-4-turbo) 72.3 56.9
ICSU (SPARQL) 72.3 -
FiDeLiS 78.3 64.3
Ours
SG-KBQA 80.3 78.2
     - Improvement +0.5% +0.5%
Table 2: F1 results (%) on WebQSP (I.I.D.) and CWQ (I.I.D.).  denotes the results we reproduced.
Competitors
On GrailQA, we compare with models that top the leaderboard [^1] https://dki-lab.github.io/GrailQA/, including RnG-KBQA (ACL 2021) (Ye et al., 2022), TIARA (EMNLP 2022) (Shu et al., 2022), DecAF (ICLR 2023) (Yu et al., 2023), Pangu (ACL 2023, SOTA before SG-KBQA, as of 19th May, 2025) (Gu et al., 2023), FC-KBQA (ACL 2023) (Zhang et al., 2023), TIARA+GAIN (EACL 2024) (Shu and Yu, 2024), and RetinaQA (ACL 2024) (Faldu et al., 2024). We also compare with few-shot LLM-based (training-free) methods: KB-BINDER (6)-R (ACL 2023) (Li et al., 2023), Pangu (Gu et al., 2023), and FlexKBQA (AAAI 2024) (Li et al., 2024). These models are SP-based. On the non-I.I.D. GrailQA, IR-based methods are uncompetitive and excluded.
On WebQSP and CWQ, we compare with IR-based models SR+NSM (ACL 2022) (Zhang et al., 2022), UNIKGQA (ICLR 2023) (Jiang et al., 2023), and EPR+NSM (WWW 2024) (Ding et al., 2024), plus LLM-based supervised fine-tuning (SFT) models including ChatKBQA (ACL 2024)(SOTA) (Luo et al., 2024a), TFS-KBQA (LREC-COLING 2024, SOTA) (Wang and Qin, 2024), and RoG (ICLR 2023) (Luo et al., 2024b). We also compare with few-shot LLM-based methods: KB-Binder (6)-R, Pangu (Codex), FlexKBQA, ToG (ICLR 2024) (Sun et al., 2024), ICSU (KSEM 2024) (Liu et al., 2024), and FiDeLiS (ACL 2025) (Sui et al., 2025). Appendix C details these models. The baseline results are collected from their papers or the GrailQA leaderboard (when available), with ToG results reproduced due to missing F1 scores on WebQSP and CWQ.
GrailQA WebQSP
Overall I.I.D. Compositional Zero-shot Overall
Model EM F1 EM F1 EM F1 EM F1 F1
SG-KBQA 85.1 88.5 93.1 94.6 78.4 83.6 84.4 87.9 80.3
   w/o SER 81.0 84.9 90.1 91.9 73.9 79.6 80.0 84.0 78.1
   w/o DED 84.3 87.8 92.6 94.0 77.1 82.4 83.7 87.2 78.2
   w/o SC 76.6 79.2 91.7 92.9 72.3 77.4 71.7 73.9 77.1
   w/o Fallback LF 81.8 84.6 92.8 94.1 77.3 81.8 78.7 81.5 78.6
Table 3: Ablation study results on the validation set of GrailQA and the test set of WebQSP.
Implementation Details
All our experiments are run on a machine with an NVDIA A100 GPU and 120 GB of RAM. For each dataset, a T5-base model is fine-tuned for 5 epochs as our logical form sketch parser. We fine-tune a LLaMA3.1-8B with LoRA Hu et al. (2022a) for 5 epochs on GrailQA and 20 epochs on WebQSP and CWQ to serve as the logical form generator. Our system parameters are selected empirically. There are only a small number of parameters to consider. As shown in the parameter study in Appendix F, our model performance shows stable patterns against the choice of parameter values. The parameter values do not take excessive fine-tuning. More implementation details are in Appendix D.
Evaluation Metrics
On GrailQA, we report the exact match (EM) and F1 scores, following the leaderboard. EM counts the percentage of test samples where the model generated logical form (an S-expression) that is semantically equivalent to the ground truth. F1 measures the answer set correctness, i.e., the F1 score of each answer set, average over all test samples. On WebQSP and CWQ, we report the F1 score as there are no ground-truth S-expressions. In addition, some baselines also report Hit, which measures whether at least one correct answer appears in the predicted answer set. Additional Hit results are reported in Appendix E.
5.2 Overall Results (Q1)
Table 1 reports the overall comparison of SG-KBQA with the baseline models on GrailQA, while Table 2 presents the corresponding results on WebQSP and CWQ. SG-KBQA shows the best results across all three datasets.
Results on GrailQA
On the overall hidden test set of GrailQA, SG-KBQA outperforms the best baseline Pangu by 4.9% and 3.3% in EM and F1 scores. While the performance improvement in the I.I.D. setting is smaller, SG-KBQA achieves substantial gains in non-I.I.D scenarios. For example, under the compositional generalization setting, it increases EM by 4.4% and F1 by 4.0% over the best baseline models. Similar performance gaps are observed under the zero-shot setting, i.e., 5.0% in EM and 2.9% in F1. Notably, under non-I.I.D. settings, the improvement in EM is consistently larger than that in F1, indicating that SG-KBQA is more capable of generating logical forms that precisely match both the questions and the KB schema, thanks to the class information that indicate the connections between the KB elements.
Most few-shot LLM-based competitors are generally not very competitive, especially under the non-I.I.D. settings and on complex questions (CWQ). A key reason is that a limited set of examples fails to sufficiently represent the diversity of KB elements necessary to address these challenges.
Results on WebQSP and CWQ
On WebQSP and CWQ, which both have I.I.D. test sets, the performance gap of different models is closer. Even in these cases, SG-KBQA still performs the best, showing its applicability. Comparing with TFS-KBQA (SOTA on WebQSP) and ChatKBQA (SOTA on CWQ), SG-KBQA improves the F1 score by 0.5% on both datasets. Among IR-based methods, UniKGQA (SOTA on WebQSP) and EPR+NSM (SOTA on CWQ) still perform much worse than SG-KBQA. The lower performance of IR-based methods is consistent with existing results (Gu et al., 2022).
@@ 5.2 Overall Results (Q1)
TIARA (T5-base) 75.3 81.9 88.4 91.2 66.4 74.8 73.3 80.7
   w SER 79.5 84.3 90.3 92.3 71.2 78.1 78.3 83.3
   w DED & SC 79.9 85.6 88.6 92.3 72.7 79.8 79.0 85.0
SG-KBQA 85.1 88.5 93.1 94.6 78.4 83.6 84.4 87.9
   w T5-base 80.6 84.9 89.9 92.6 73.8 81.0 79.4 83.3
Table 4: Module applicability results on the validation set of GrailQA.